OpenStudy (anonymous):

Equilateral triangle has sides: 3x-3; 5x-2y; 6y-4. Find values of x and y. I keep getting crazy answers when I substitute back!

OpenStudy (anonymous):

3x-3=5x-2y=6y-4

OpenStudy (anonymous):

5 x - 2 y = 6 y - 4 Solve the above in terms of y, you will get \[ y=\frac{1}{8} (5 x+4 \] replace y in 3x-3=5x-2y and solve for x. Let me know what you find for x and then for y

OpenStudy (anonymous):

\[y=\frac{1}{8} (5 x+4) \] in my post above

OpenStudy (mertsj):

I got x = -8/3 and y = -7/6 What did you get?

OpenStudy (anonymous):

That is what I got.

OpenStudy (mertsj):

I guess the asker is no longer interested.

OpenStudy (anonymous):

Hi you all... sorry was busy elsewhere. i want to figure out again now that i know your answers.Thks for help. will come in again later!!

OpenStudy (anonymous):

Hum, negative side......

OpenStudy (anonymous):

fter lots of stupid mistakes, I finally got the same. Now I must check if the lengths of the sides of the triangle all make sense. Thanks!